# FreezeAsGuard: Mitigating Illegal Adaptation of Diffusion Models via Selective Tensor Freezing

### v1

@@ Abstract
Text-to-image diffusion models can be fine-tuned in custom domains to adapt to specific user preferences, but such unconstrained adaptability has also been utilized for illegal purposes, such as forging public figures’ portraits and duplicating copyrighted artworks. Most existing work focuses on detecting the illegally generated contents, but cannot prevent or mitigate illegal adaptations of diffusion models. Other schemes of model unlearning and reinitialization, similarly, cannot prevent users from relearning the knowledge of illegal model adaptation with custom data. In this paper, we present FreezeAsGuard, a new technique that addresses these limitations and enables irreversible mitigation of illegal adaptations of diffusion models. The basic approach is that the model publisher selectively freezes tensors in pre-trained diffusion models that are critical to illegal model adaptations, to mitigate the fine-tuned model’s representation power in illegal domains but minimize the impact on legal model adaptations in other domains. Such tensor freezing can be enforced via APIs provided by the model publisher for fine-tuning, can motivate users’ adoption due to its computational savings. Experiment results with datasets in multiple domains show that FreezeAsGuard provides stronger power in mitigating illegal model adaptations of generating fake public figures’ portraits, while having the minimum impact on model adaptation in other legal domains. The source code is available at: https://github.com/pittisl/FreezeAsGuard.
@@ 1 Introduction
Text-to-image diffusion models [43; 42] are powerful tools to generate high-quality images aligned with user prompts. After being pre-trained by model publishers to embed world knowledge from large image data [48], open-sourced diffusion models, such as Stable Diffusion (SD) [8; 9], can be conveniently [^1] Many APIs, such as HuggingFace Diffusers [52], can be used for fine-tuning with the minimum user efforts. adapted by users to generate their preferred images, through fine-tuning with custom data in specific domains. For example, diffusion models can be fine-tuned on cartoon datasets to synthesize avatars in video games [45], or on datasets of landscape photos to generate wallpapers [11].
An increasing risk of democratizing open-sourced diffusion models, however, is that the capability of model adaptation has been utilized for illegal purposes, such as forging public figures’ portraits [24; 26], duplicating copyrighted artworks [28], and generating sexual content [27]. Most existing efforts aim to deter attempts of illegal model adaptation with copyright detection [58; 18; 19], which embeds invisible but detectable watermarks into training data and further generated images, as shown in Figure 1. However, such detection only applies to misuse of training data, and does not mitigate the user’s capability of illegal model adaptation.
Instead, an intuitive approach to mitigation is content filtering. However, filtering user prompts [20] can be bypassed by fine-tuning the model to align innocent prompts with illegal image contents [51], and filtering the generated images [6] is often overpowered with high false-positive rates [1]. Data poisoning techniques can avoid false positives by injecting invisible perturbations into training data [54; 57; 49], but cannot apply when public data or users’ private data is used for fine-tuning. Recent unlearning methods allow model publishers to remove knowledge needed for illegal adaptation by modifying model weights [21; 25; 53; 59] , but cannot prevent relearning such knowledge via fine-tuning.
The key limitation of the aforementioned existing techniques is that they focus on modifying the training data or model weights, but such modification can be easily reversed by users via fine-tuning with their custom data. Such modification, furthermore, cannot precisely focus the mitigation power only in illegal domains without affecting model adaptation in other innocent domains, due to the high ambiguity when fine-tuning diffusion models in different domains.
To prevent users from reversing the mitigation maneuvers being applied on diffusion models, in this paper we present a new technique, namely FreezeAsGuard, which advocates a fundamental shift that constrains the trainability of diffusion model’s tensors in fine-tuning. As shown in Figure 1, the model publisher selectively freezes tensors in pre-trained models that are critical to the convergence of fine-tuning in illegal domains, to limit the model’s representation power of being fine-tuned in illegal domains. In practice, since most users fine-tune diffusion models by following instructions and using APIs provided by model publishers, tensor freezing can be effectively enforced by model publishers through these APIs. Meanwhile, since freezing tensors significantly reduces the computational costs of fine-tuning, users are well motivated to adopt tensor freezing in fine-tuning practices.
@@
Figure 2: FreezeAsGuard ensures that portrait images generated by fine-tuned model in illegal domains cannot be recognizable as target subjects. In contrast, unlearning schemes (UCE [25] and IMMA [59]) cannot prevent the unlearned knowledge of illegal domains from being relearned in fine-tuning.
Figure 3: Mask learning and fine-tuning are combined as bilevel optimization
@@
The major challenge is how to properly evaluate the importance of tensors in model fine-tuning. Popular attribution-based importance metrics [37; 40] are mainly used in model pruning with fixed weight values, but cannot reflect the impact of weight variations in fine-tuning. Such impact of weight variations, in fact, cannot be condensed into a single importance metric, due to the randomness and interdependencies of weight updates in fine-tuning iterations. Instead, as shown in Figure 3, we formulate the selection of frozen tensors as a trainable binary mask. Given a required ratio of frozen tensors specified by the model publisher, we optimize such selection with training data samples in illegal domains, through bilevel optimization that combines the iterative process of mask learning and the iterations of model fine-tuning. In this way, we ensure that the mask being trained can timely learn the impact of weight variations on the training loss during fine-tuning.
@@
With frozen tensors, the model’s representation power should be retained when fine-tuned in innocent domains. Hence, we further incorporate training samples from innocent domains into the bilevel optimization, to provide suppressing signals for selecting tensors being frozen. Hence, the learned mask of freezing tensors should skip tensors that are important to fine-tuning in innocent domains.
We evaluated FreezeAsGuard in a challenging mitigation of illegal model adaptation: generating fake portraits of public figures, with open-sourced SD models. We use a self-collected dataset of 25 public figures’ portraits as illegal domain, Modern-Logo-v4 [4] and H&M-Clothes [3] datasets as innocent domain, and competitive model unlearning schemes as baselines. Our main findings are as follows:
@@
  • FreezeAsGuard has strong mitigation power in illegal domains. It reduces the quality of images generated by fine-tuned model by 14% compared to baselines, and ensures the generated images to be unrecognizable as target subjects, as shown in Figure 2 and Section 4.2.
@@
  • FreezeAsGuard has the minimum impact on legal modal adaptation in innocent domains. It can achieve on-par quality of the generated images with regular full fine-tuning on innocent datasets. Compared to the competitive baselines, it improves the accuracy by up to 8%.
@@
  • FreezeAsGuard has high compute efficiency. Compared to baseline schemes, it can save up to 48% GPU memory and 21% wall-clock time of model fine-tuning for innocent users.
@@ 2 Background & Motivation
2.1 Fine-tuning Diffusion Models in illegal domains
Given text prompts $y$ and images $x$ as training data, fine-tuning a diffusion model approximates the conditional distribution $p(x|y)$ by learning to reconstruct images that are progressively blurred with Gaussian noise $\epsilon$ over step $t=1,...,T$. The training objective is to minimize the reconstruction loss:
$\mathcal{L}_{DM}=\mathbb{E}_{x,y,\epsilon\sim\mathcal{N}(0,1),t}\left[\|\epsilon-\epsilon_{\theta}(\mathcal{E}(x_{t}),t,\tau(y))\|^{2}_{2}\right],$ (1)
where $\mathcal{E}(\cdot)$ is the encoder of a pretrained VAE, $\tau(\cdot)$ is a pretrained text encoder, and $\epsilon_{\theta}(\cdot)$ is a denoising model with trainable parameters $\theta$. Most existing diffusion models adopt the UNet architecture [44] for the denoising model, which is then used to generate images from noise, conditioned by user prompts.
In fine-tuning, the diffusion model learns new knowledge in illegal domains by adopting generic knowledge in the pre-trained model [14]. For example, knowledge about “a green beetle” in illegal domains can be a combination of knowledge on “hornet” and “emerald” in pre-trained model. This behavior implies that fine-tuning in different domains shares the same knowledge base, and it is hence challenging to focus the mitigation power in illegal domains without affecting fine-tuning in innocent domains. This challenge motivates us to regulate FreezeAsGuard’s mitigation power by incorporating training samples in innocent domains, when selecting tensors being frozen for illegal domains.
In most fine-tuning practices, LoRA adapters [32; 46] and prompt tuning [23] are used to reduce memory costs. These methods, however, restrict the number of trainable UNet parameters and reduce the representation power of fine-tuned models in illegal domains. In this paper, we instead consider a more challenging scenario of tensor freezing, where all UNet parameters are trainable in fine-tuning.
@@ 2.1 Fine-tuning Diffusion Models in illegal domains
Table 1: Quality of generated images with different model compoents being frozen when fine-tuning SD v1.5 [8] on the captioned pokemon dataset [5], using CLIP [29, TOPIQ [15], and FID [30] image quality metrics
Figure 4: Generated images with different model components being frozen, when using the prompt “a cartoon pikachu with a pink dress and a pink bow”
@@ 2.2 Partial Model Fine-tuning
Different model components have specialized functionalities, and hence fine-tuning only some of them results in different model performances. For diffusion models, we cannot mitigate illegal model adaptations in illegal domains by simply excluding some layers from being fine-tuned, because shallow layers provide primary image features and deep layers enforce domain-specific semantics [56]. They are, hence, both essential to the performance of the fine-tuned models in innocent domains.
Similarly, although large diffusion models have modularized structures, it is inappropriate to completely exclude specific model components from being fine-tuned. As shown in Table 1 and Figure 4, freezing critical components such as attention projectors and time embeddings can cause non-negligible quality drop in generated images. Even when freezing the same amount of model weights (e.g., random 50%), the exact distribution of frozen weights could also affect the generated images’ quality. Such heterogeneity motivates us to instead seek for globally optimal selections of freezing tensors across all model components, by jointly taking all model components into bilevel optimization.
@@ 3 Method
Our design of FreezeAsGuard builds on bilevel optimization, which embeds one optimization problem within another [16; 39; 22]. This bilevel optimization can be formulated as
@@
$\mathbf{m}^{*}$ $=\arg\min_{\mathbf{m}}\left(-\mathcal{L}_{\boldsymbol{\theta}^{*}(\mathbf{m})}(\mathbf{x}_{illegal}),\mathcal{L}_{\boldsymbol{\theta}^{*}(\mathbf{m})}(\mathbf{x}_{innocent})\right)$ (2)
[…]
@@
where $\mathbf{m}$ is the binary mask of selecting frozen tensors, $\mathbf{m}^{*}$ is the optimized binary mask, $\boldsymbol{\theta}(\mathbf{m})$ represents the model tensors frozen by $\mathbf{m}$, and $\boldsymbol{\theta}^{*}(\mathbf{m})$ is the converged $\boldsymbol{\theta}(\mathbf{m})$ after fine-tuning. $\mathbf{x}_{illegal}$ and $\mathbf{x}_{innocent}$ denote training samples in illegal domains ($\mathcal{D}_{illegal}$) and innocent domains ($\mathcal{D}_{innocent}$), respectively. Such bilevel optimization is illustrated in Figure 5. The lower-level problem in Eq. (3) is a simulated user loop that the user fine-tunes the diffusion model towards convergence by minimizing the loss over both illegal and innocent domains. The upper-level problem in Eq. (2) is a mask learning loop that the model publisher learns $\mathbf{m}$ to mitigate the diffusion model’s representation power when being fine-tuned in illegal domains, without affecting fine-tuning in innocent domains.
Some existing work adopts differentiable image quality metrics (e.g., MS-SSIM [50]) to measure the convergence of fine-tuning, but suffers from gradient explosion during optimization [54]. Instead, in FreezeAsGuard we use the standard diffusion loss in Eq. (1), and adopt tensor-level freezing to ensure sufficient granularity [^2] Most existing diffusion models have parameter sizes between 1B and 3.5B, which correspond to at least 686 tensors over the UNet-based denoiser., without incurring extra computing costs.
@@
To apply the gradient solver, $\mathbf{m}$ and $\boldsymbol{\theta}(\mathbf{m})$ should have differentiable dependencies with the loss function calculation. We model $\boldsymbol{\theta}(\mathbf{m})$ through the weighted summation of pre-trained model tensors $\boldsymbol{\theta}_{pre}$ and fine-tuned model tensors $\boldsymbol{\theta}_{ft}$, such that
$\boldsymbol{\theta}(\mathbf{m})=\mathbf{m}\odot\boldsymbol{\theta}_{pre}+(\mathbf{1}-\mathbf{m})\odot\boldsymbol{\theta}_{ft},\vspace{-0.05in}$ (4)
@@
where $\odot$ denotes element-wise multiplication. From the user’s perspective, fine-tuning the partially frozen model $\boldsymbol{\theta}(\mathbf{m})$ is equivalent to fine-tuning the $\boldsymbol{\theta}_{ft}$ component. In that case, $\boldsymbol{\theta}_{ft}$ is controlled by the constraint in Eq. (3). To improve compute efficiency, we initialize $\boldsymbol{\theta}_{ft}$ as the fully fine-tuned model tensors in both illegal and innocent domains, and gradually enlarge the scope of tensors being frozen. Besides, since $\mathbf{m}$ is discrete and not differentiable, in practice we adopt a continuous form $\mathbf{m}(\mathbf{w})=\sigma(\mathbf{w}/T)$ that applies sigmoid function $\sigma(\cdot)$ over a trainable tensor $\mathbf{w}$. We also developed code optimizations to allow vectorized gradient calculations, and details can be found in Appendix A.

### v2

@@ Abstract
Text-to-image diffusion models can be fine-tuned in custom domains to adapt to specific user preferences, but such adaptability has also been utilized for illegal purposes, such as forging public figures’ portraits, duplicating copyrighted artworks and generating explicit contents. Existing work focused on detecting the illegally generated contents, but cannot prevent or mitigate illegal adaptations of diffusion models. Other schemes of model unlearning and reinitialization, similarly, cannot prevent users from relearning the knowledge of illegal model adaptation with custom data. In this paper, we present FreezeAsGuard, a new technique that addresses these limitations and enables irreversible mitigation of illegal adaptations of diffusion models. Our approach is that the model publisher selectively freezes tensors in pre-trained diffusion models that are critical to illegal model adaptations, to mitigate the fine-tuned model’s representation power in illegal adaptations, but minimize the impact on other legal adaptations. Experiment results in multiple text-to-image application domains show that FreezeAsGuard provides 37% stronger power in mitigating illegal model adaptations compared to competitive baselines, while incurring less than 5% impact on legal model adaptations. The source code is available at: https://github.com/pittisl/FreezeAsGuard.
@@ 1 Introduction
Text-to-image diffusion models [44; 43] are powerful tools to generate high-quality images aligned with user prompts. After pre-trained by model publishers to embed world knowledge from large image data [49], open-sourced diffusion models, such as Stable Diffusion (SD) [9; 10], can be conveniently adapted by users to generate their preferred images [^1] Many APIs, such as HuggingFace Diffusers [56], can be used for fine-tuning open-sourced diffusion models with the minimum user efforts., through fine-tuning with custom data in specific domains. For example, diffusion models can be fine-tuned on cartoon datasets to synthesize avatars in video games [46], or on datasets of landscape photos to generate wallpapers [11].
@@
Figure 2: FreezeAsGuard ensures that portraits (left) and artworks (right) generated by diffusion models in illegal classes cannot be recognizable as target objects, even if the model has been fine-tuned with data samples in illegal classes. In contrast, unlearning schemes (UCE [23] and IMMA [65]) cannot prevent the unlearned knowledge of illegal classes from being relearned in fine-tuning.
An increasing risk of democratizing open-sourced diffusion models, however, is that the capability of model adaptation has been utilized for illegal purposes, such as forging public figures’ portraits [22; 24], duplicating copyrighted artworks [26], and generating explicit content [25]. Most existing efforts aim to deter attempts of illegal model adaptation with copyright detection [64; 16; 17], which embeds invisible but detectable watermarks into training data and further generated images, as shown in Figure 1. However, such detection only applies to misuse of training data, and does not mitigate the user’s capability of illegal model adaptation. Users can easily bypass such detection by collecting and using their own training data without being watermarked (e.g., users’ self-taken photos of public figures).
Instead, an intuitive approach to mitigation is content filtering. However, filtering user prompts [19] can be bypassed by fine-tuning the model to align innocent prompts with illegal image contents [55], and filtering the generated images [7] is often overpowered with high false-positive rates [3]. Data poisoning techniques can avoid false positives by injecting invisible perturbations into training data [59; 62; 51], but cannot apply when public web data or users’ private data is used for fine-tuning. Recent unlearning methods allow model publishers to remove knowledge needed for illegal adaptation by modifying model weights [20; 23; 57; 65] , but cannot prevent relearning such knowledge via fine-tuning.
The key limitation of these techniques is that they focus on modifying the training data or model weights, but such modification can be reversed by users via fine-tuning with their own data. Such modification, further, cannot restrain the mitigation power only in illegal data classes (e.g., public figures’ portraits) without affecting model adaptation in other legal data classes (e.g., the user’s own portraits), due to the high ambiguity and possible overlap between these classes.
To prevent users from reversing the mitigation maneuvers being applied, in this paper we present FreezeAsGuard, a new technique that constrains the trainability of diffusion model’s tensors in fine-tuning. As shown in Figure 1, the model publisher selectively freezes tensors in pre-trained models that are critical to fine-tuning in illegal classes (e.g., public figures’ portraits), to limit the model’s representation power of being fine-tuned in illegal classes. In practice, since most illegal users are not professional and fine-tune diffusion models by simply following the instructions provided by model publishers, tensor freezing can be effectively enforced by model publishers through these instructions, to guide the users to adopt tensor freezing. Essentially, since freezing tensors lowers the trainable model parameters and reduces the computing costs of fine-tuning, users would be well motivated to adopt tensor freezing in fine-tuning practices.
Figure 3: Mask learning and fine-tuning as a bilevel optimization
@@
The major challenge is how to properly evaluate the importance of tensors in model fine-tuning. Popular attribution-based importance metrics [38; 41] are used in model pruning with fixed weight values, but cannot reflect the impact of weight variations in fine-tuning. Such impact of weight variations, in fact, cannot be condensed into a single importance metric, due to the randomness and interdependencies of weight updates in fine-tuning iterations.
Instead, as shown in Figure 3, we formulate the selection of frozen tensors in all the illegal classes as one trainable binary mask. Given a required ratio of frozen tensors specified by model publisher, we optimize such selection with training data in all the involved illegal classes, through bilevel optimization that combines the iterative process of mask learning and iterations of model fine-tuning. In this way, the mask being trained can timely learn the impact of weight variations on the training loss during fine-tuning.
@@
With frozen tensors, the model’s representation power should be retained when fine-tuned on other legal classes (e.g., user’s own portraits). Hence, we incorporate training samples from legal classes into the bilevel optimization, to provide suppressing signals for selecting tensors being frozen. Hence, the learned mask of freezing tensors should skip tensors that are important to fine-tuning in legal classes.
We evaluated FreezeAsGuard in three different domains of illegal model adaptations: 1) forging public figures’ portraits, 2) duplicating copyrighted artworks and 3) generating explicit contents. For each domain, we use open-sourced or self-collected datasets, and randomly select different data classes as illegal and legal classes. We use competitive model unlearning schemes as baselines, and multiple metrics to measure image quality. Our findings are as follows:
@@
  • FreezeAsGuard has strong mitigation power in illegal classes. Compared to the competitive baselines, it further reduces the quality of images generated by fine-tuned model by up to 37%, and ensures the generated images to be unrecognizable as subjects in illegal classes.
@@
  • FreezeAsGuard has the minimum impact on modal adaptation in legal classes. It ensures on-par quality of the generated images compared to regular full fine-tuning on legal data, with a difference of at most 5%.
@@
  • FreezeAsGuard has high compute efficiency. Compared to full fine-tuning, it can save up to 48% GPU memory and 21% wall-clock computing time.
@@ 2 Background & Motivation
2.1 Fine-Tuning Diffusion Models
Given text prompts $y$ and images $x$ as training data, fine-tuning a diffusion model approximates the conditional distribution $p(x|y)$ by learning to reconstruct images that are progressively blurred with noise $\epsilon$ over step $t=1,...,T$. Training objective is to minimize the reconstruction loss:
$\mathcal{L}_{\theta}=\mathbb{E}_{x,y,\epsilon\sim\mathcal{N}(0,1),t}\left[\|\epsilon-\epsilon_{\theta}(\mathcal{E}(x_{t}),t,\tau(y))\|^{2}_{2}\right],\vspace{-0.1in}$ (1)
where $\mathcal{E}(\cdot)$ is the encoder of a pretrained VAE, $\tau(\cdot)$ is a pretrained text encoder, and $\epsilon_{\theta}(\cdot)$ is a denoising model with trainable parameters $\theta$. Most diffusion models adopt UNet architecture [45] as the denoising model.
In fine-tuning, the diffusion model learns new knowledge by adapting the generic knowledge in the pre-trained model [13]. For example, new knowledge about “a green beetle” can be a combination of generic knowledge on “hornet” and “emerald”. This behavior implies that fine-tuning in different classes may share the same knowledge base, and it is challenging to focus the mitigation power in illegal classes without affecting fine-tuning in other legal classes. This challenge motivates us to regulate FreezeAsGuard’s mitigation power by incorporating training samples in legal classes, when selecting tensors being frozen for illegal classes.
@@ 2.1 Fine-Tuning Diffusion Models
Table 1: Quality of generated images with different model compoents being frozen, using CLIP [27], TOPIQ [14], and FID [28] image quality metrics and the captioned pokemon dataset [6]
Figure 4: Generated images with different model components being frozen, with prompt “a pikachu with a pink dress and a pink bow”
@@ 2.2 Partial Model Fine-tuning
An intuitive solution to mitigating illegal model adaptation is to only allow fine-tuning some layers or components of the diffusion model. However, this solution is ineffective in practice, because shallow layers provide primary image features and deep layers enforce domain-specific semantics [61]. They are, hence, both essential to the performance of the fine-tuned models in legal classes. Similarly, as shown in Table 1 and Figure 4, freezing critical model components such as attention projectors and time embeddings can cause large quality drop in generated images. Even when freezing the same amount of model weights (e.g., random 50%), the exact distribution of frozen weights could also affect the generated images’ quality. Such heterogeneity motivates us to instead seek for globally optimal selections of freezing tensors across all model components, by jointly taking all model components into bilevel optimization.
@@ 3 Method
Our design of FreezeAsGuard builds on bilevel optimization, which embeds one optimization problem within another and both of them are multi-objective optimizations [15; 40; 21]. This bilevel optimization can be formulated as
@@
$\mathbf{m}^{*}$ $=\arg\min_{\mathbf{m}}\left(-\mathcal{L}_{\boldsymbol{\theta}^{*}(\mathbf{m})}(\mathbf{x}_{illegal}),\mathcal{L}_{\boldsymbol{\theta}^{*}(\mathbf{m})}(\mathbf{x}_{legal})\right)$ (2)
s.t. $\ \ \ \boldsymbol{\theta}^{*}(\mathbf{m})=\arg\min_{\boldsymbol{\theta}(\mathbf{m})}\left(\mathcal{L}_{\boldsymbol{\theta}(\mathbf{m})}(\mathbf{x}_{illegal}),\mathcal{L}_{\boldsymbol{\theta}(\mathbf{m})}(\mathbf{x}_{legal})\right),$ (3)
@@
where $\mathbf{m}$ is the binary mask of selecting frozen tensors, $\mathbf{m}^{*}$ is the optimized binary mask, $\boldsymbol{\theta}(\mathbf{m})$ represents the model tensors frozen by $\mathbf{m}$, and $\boldsymbol{\theta}^{*}(\mathbf{m})$ is the converged $\boldsymbol{\theta}(\mathbf{m})$ after fine-tuning. $\mathbf{x}_{illegal}$ and $\mathbf{x}_{legal}$ denote training samples in all the illegal classes ($\mathcal{C}_{illegal}$) and legal classes ($\mathcal{C}_{legal}$), respectively. Such bilevel optimization is illustrated in Figure 5. The lower-level problem in Eq. (3) is a simulated user loop that the user fine-tunes the diffusion model by minimizing the loss over both illegal and legal classes. The upper-level problem in Eq. (2) is a mask learning loop that learns $\mathbf{m}$ to mitigate the model’s representation power when fine-tuned in illegal classes, without affecting fine-tuning in legal classes. We use the standard diffusion loss in Eq. (1) and adopt tensor-level freezing to ensure sufficient granularity [^2] Most existing diffusion models have parameter sizes between 1B and 3.5B, which correspond to at least 686 tensors over the UNet-based denoiser., without incurring extra computing costs.
@@
To apply the gradient solver, $\mathbf{m}$ and $\boldsymbol{\theta}(\mathbf{m})$ should have differentiable dependencies with the loss function. We model $\boldsymbol{\theta}(\mathbf{m})$ through the weighted summation of pre-trained model tensors $\boldsymbol{\theta}_{pre}$ and fine-tuned model tensors $\boldsymbol{\theta}_{ft}$, such that
$\boldsymbol{\theta}(\mathbf{m})=\mathbf{m}\odot\boldsymbol{\theta}_{pre}+(\mathbf{1}-\mathbf{m})\odot\boldsymbol{\theta}_{ft},$ (4)
@@
where $\odot$ denotes element-wise multiplication. From the user’s perspective, fine-tuning the partially frozen model $\boldsymbol{\theta}(\mathbf{m})$ is equivalent to fine-tuning $\boldsymbol{\theta}_{ft}$, controlled by Eq. (3). To improve compute efficiency, we initialize $\boldsymbol{\theta}_{ft}$ as the fully fine-tuned model tensors on both illegal and legal classes, and gradually enlarge the scope of tensor freezing. Since $\mathbf{m}$ is discrete and not differentiable, we adopt a continuous form $\mathbf{m}(\mathbf{w})=\sigma(\mathbf{w}/T)$ that applies sigmoid function $\sigma(\cdot)$ over a trainable tensor $\mathbf{w}$. We also did code optimizations for vectorized gradient calculations as in Appendix A.
Note that, although we made $\mathbf{m}$ differentiable in bilevel optimizations, the optimized values in $\mathbf{m}^{*}$ will be rounded to binary, to ensure complete freezing of selected tensors.